The common stock for the Bestsold Corporation sells for $62. If a new issue is​ sold, the flotation costs are estimated to be 9 percent. The company pays 60 percent of its earnings in​ dividends, and a $3.60 dividend was recently paid. Earnings per share 5 years ago were $4.00. Earnings are expected to continue to grow at the same annual rate in the future as during the past 5 years. The​ firm's marginal tax rate is 32 percent. Calculate the cost of:

​(a​) internal common equity and

​(b​) external common equity.
Answer #1

EPS0 = 3.60/0.60 = $6.00

g = (6/4)1/5 - 1 = 8.45%

a.

For Internal Common Equity,

Using Constant Growth Model,

r = 3.60(1.0845)/62 + 0.0845

r = 14.75%

b.

For External Common Equity,

Using Constant Growth Model,

r = 3.60(1.0845)/(0.91*62) + 0.0845

r = 15.37%